OpenStudy (anonymous):

Sum of: (-1)^(n+1)*n^(-2)*n^(-n+4)*(1+3/n)^(n^2) with n=1 to +inifinity. Need to tell if the series converges or diverges, don't need to know the specific number it goes to. I've done the root test on it but got stuck when it came to dealing with the exponents. Please help, thanks

OpenStudy (tkhunny):

\((-1)^{n+1}\) -- Okay, it alternates, so we need to establish only that the terms converge to zero. Then we're done. Well, assuming it converges. \(\dfrac{\left(1+\dfrac{3}{n}\right)^{n^2}}{n^{2}\cdot n^{n-4}}\) Hmmm... What do you think? \(\lim\limits_{x\rightarrow\infty}\left(1+\dfrac{3}{n}\right)^{n} = e^{3}\). Does that get us anywhere?

OpenStudy (anonymous):

So that limit you show at the bottom of your response is a rule right? So e^3 would be the numerator, then you can combine the denominator

OpenStudy (anonymous):

After you rewrite the equation, is it then that you do the root test?

OpenStudy (tkhunny):

No, e^3 is not the numerator. There's an n^2 up there, not just an n.

OpenStudy (anonymous):

I did the root test and it brought the n^2 down to an n

OpenStudy (tkhunny):

Maybe compare to \(\dfrac{e^{3n}}{n^{n-2}}\). Do we get l'Hospital?

OpenStudy (anonymous):

We definitely get 1^infinity for the numerator, and 1 for the denominator

OpenStudy (tkhunny):

That's no good. They are \(\dfrac{\infty}{\infty}\). Can we use it?

OpenStudy (anonymous):

i thought l'hopitals rule could also be 1^infinity?

OpenStudy (anonymous):

At least my professor says it can...

OpenStudy (tkhunny):

It can, but that's not what we have.

OpenStudy (anonymous):

okay, so infinity/infinity, i figured out why i thought the other way

OpenStudy (anonymous):

if you have \[\frac{ \infty }{\infty}\] then you would take the derivative of each the top and bottom

OpenStudy (tkhunny):

Yeah, no. That won't get us anywhere. We need to rewrite to understand. \(\left(\dfrac{\left(1+\dfrac{3}{n}\right)^{n}}{n}\right)^{n}\cdot n^{2} < \left(\dfrac{e}{n}\right)^{n}\cdot n^{2} = \dfrac{e^{n}}{n^{n-2}}\) and that clearly converges to zero after n > 2.